STATEMENT OF FINANCIAL POSITION


Accounts Payable ₱ 600,000

Unearned Revenue 1,210,000

Salaries Payable- current 120,000

Owner’s Withdrawals 280,000

Owner’s Capital 1,500,000

Comprehensive Income 3,000,000

Total Current Assets 80% of total liabilities

Cash 250,000

Inventories 480,000

Prepaid Expenses 600,000

Property, Plant and Equipment 5,166,000

Accounts Receivable 214,000


Look for:

A. CURRENT RATION

B. QUICK RATIO

C. CURRENT ASSETS TO WORKING CAPITAL

D. DEFENSIVE INTERVAL RATIO

E. DEBT TO EQUITY RATIO

F. EQUITY TO DEBT RATIO

G. DEBT RATIO

H. EQUITY RATIO

1
Expert's answer
2021-10-27T18:56:03-0400

A.

"Current Ratio = \frac{Current assets}{Current liabilities}"

Where;

Current assets = cash + inventories + prepaid expenses + accounts Receivable

= 250000 + 480000 + 600000 + 214000 = 1,544,000

Current liabilities = accounts payable + unearned revenue + salaries payable

= 600000 + 1210000 + 120000 = 1,930,000

"Current Ratio = \frac{1,544,000}{1,930,000}=0.80"


B.

"Quick Ratio = \frac{Current assets-inventories}{Current liabilities}"


"QuickRatio = \frac{1,544,000-480,000}{1,930,000}=0.55"


C.

"Current assets to working capital = \frac{Current assets}{Working capital}"


"Current assets to working capital = \frac{1,544,000}{1,544,000-1,930,000}=-4.0"


D.

Cash operating expenses are not given, so it is not possible to calculate defensive interval ratio.


E.

"Debt-to-equity =\frac{Total liabilities}{Total equity}"

Where;

Total equity = Owner’s Withdrawals + Owner’s Capital + Comprehensive Income

= 280,000 + 1,500,000 + 3,000,000 = 4,780,000

Total liabilities = 1,930,000

"Debt-to-equity =\frac{1,930,000}{4,780,000}=0.40"


F.

"Equity-to-debt=\frac{Total equity}{Total liabilities}"


"Equity-to-debt=\frac{4,780,000}{1,930,000}=2.5"


G.

"Debt ratio=\frac{Total liabilities}{Total assets}"

Where;

Total assets = 1544000 + 5166000 = 6,710,000

"Debt ratio=\frac{1,930,000}{6,710,000}=0.29"


H.

"Equityratio=\frac{Total equity}{Total assets}"


"Equityratio=\frac{4,780,000}{6,710,000}=0.71"